Please help me with this word problem.One day, the temperature at 2:00 A.M. was 43°F, and by 3:00 P.M. the temperature was 56°F. What was the hourly rate of temperature change? I came up with 1°F/h. Thank-you

Answer by stanbon(75887)

One day, the temperature at 2:00 A.M. was 43°F, and by 3:00 P.M. the temperature was 56°F. What was the hourly rate of temperature change? I came up with 1°F/h.
-----------------
The hourly rate is the slope. You have two points(2,43) and (15,56)
slope = (56-43)/(15-2) = 13/13 = 1 degree per hour on the average.
======
Your answer looks good.
=========
Cheers,
Stan H.